Factor completely -9 + 18x^2 + 21x Please show step by step the factoring process, thank you
Answer by edjones(8007)

18x^2+21x-9
=3(6x^2+7x-3)
6*-3=-18; Two factors of -18 whose sum is 7 are 9 and -2.
3(6x^2+9x-2x-3)
=3(3x(2x+3)-(2x+3))
=3(3x-1)(2x+3)
.
Ed